Sandhill Corporation has fixed costs of $323,500. It has a unit selling price of $6.35, unit variable cost of $4.85, and a target net income of $1,520,000.

Compute the required sales in units to achieve its target net income.

Required sales

enter the required sales in units

units
Answer #1

Required sales in units = (fixed cost + targeted net income)/ contribution margin per unit

Contribution margin per unit = unit selling price – unit variable cost = 6.35-4.85 = $1.50

Required sales in units = (323500+1520000)/1.50 = 1229000 units